Suppose you have been hired as a management consultant by a major oil company to help it optimally price gasoline at its service stations. Your client wants to know what wl happen to gasoline demand if it increases gasoline prices by one cent higher than its nearest competitors. One of the members of your consulting team, Debbie, shares that one time in college she stopped buying gasoline from a service station that was one cent more expensive Based on this story, should you conclude that demand will fall to zero if the client raises gas prices by one cent? since this is an which can lead to Instead of relying on Debbies college story to make a conclusion, you decide it is smarter to collect and analyze, O A. a survey of anecdotes collected on the Internet. B. a survey of other team members anecdotes. C. a large amount of empirical data. D. a small sample of empirical data. You are able to collect the following data on the impact of a one-cent increase in gasoline prices. Calculate the percentage change in demand for each service station. (Round your responses to two decimal places. If the change is negative, be sure to include a minus sign.) Service Station Gasoline Sold (thousands of gallons) Gasoline Sold (thousands of gallons) Percentage Change With No Price Increase 120 160 With 1-Cent Price Increase 90 140
Answer #1

NO​, since this is an ARGUMENT BY ANECDOTE​  which can lead to WRONG CONCLUSIONS.

a large amount of empirical data
Instead of relying on​ Debbie's college story to make a​ conclusion, you decide it is smarter to collect and analyze a large amount of empirical data

Percentage change

A

90-120/120*100= -30/120*100= -1/4*100= -25%

B

140-160/160*100= -20/160*100=-1/8*100= -12.5%